Two balls are drawn in succession without replacement from a box containing 4 yellow balls and 3 white balls. Let Z be the random variable representing the number of white balls. Find the values of the random variable Z.




1
Expert's answer
2022-02-21T15:56:46-0500

"P(Z=0)=\frac{4}{7}*\frac{3}{6}=\frac{2}{7}."

"P(Z=1)=\frac{3}{7}*\frac{4}{6}*2=\frac{4}{7}."

"P(Z=1)=\frac{3}{7}*\frac{2}{6}=\frac{1}{7}."